Cheryl has invested $600 in a mutual fund that charges 6% front-end load. She plans to keep her investment in the fund for the next 10 years. What will be her average annual return, if she expects the fund to earn a 5% reported return in an average year? 4.35% 4.7% -1.3% 5%
Answer #1

SOLUTION :


Upload front = 6% = 0.06 in decimals.


So, effective cost of $1  after front-end-upload of $0.06 = 1 + 0.06 = 1.06 ($)


Average return in a year = 5% = $0.05 per dollar of cost.


So, 


Effective return per dollar when effective cost is $1.06

= 0.05 / 1.06

= 0.047

= 4.7 % : 2nd Option : ANSWER.

answered by: Tulsiram Garg
Answer #2

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASEE AutoSum7 B a copy 2 Clear Sort & Filter Editing Find & Select Cells 740 PE EPS, STOCK DIVIDEND - Microsoft Excel (Product A